# Fast Scrambling in Hyperbolic Ising Model

### v1

@@ Abstract
I Introduction
Quantum many-body chaos and scrambling have attracted a great deal of attention in recent years, from more theoretical endeavors like studying the scrambling dynamics of black holes using their dual descriptions Kobrin et al. (2021); Fu and Sachdev (2016); Asaduzzaman et al. (2024a); Brown et al. (2023); Schuster et al. (2022); Cotler et al. (2017); Joshi et al. (2022) to more practical applications like investigating scrambling in chemical reactions to quantum cryptography Chamon et al. (2024, 2022) and many more.
In almost all of these studies, the Sachdev-Ye-Kitaev (SYK) model has been the center of attention as the quintessential example of a model that exhibits fast scrambling and quantum chaos Sachdev and Ye (1993); noa (a, b); Maldacena and Stanford (2016). However, due to the all-to-all and random interactions in this model, any kind of classical and quantum simulation can be very challenging and often requires various truncations and approximations Xu et al. (2020); García-García et al. (2021); Cao et al. (2020); Sünderhauf et al. (2019). In this letter, we advocate for a new model that is also maximally chaotic and exhibits fast scrambling but with only local site-dependent interactions.  [^1] There are other examples of fast scrambling models besides the SYK model but to the best of our knowledge they all require longer-range interactions than nearest neighbor interactions to be present in the model Hashizume et al. (2021); Li et al. (2020); Lashkari et al. (2013); Belyansky et al. (2020); Yin and Lucas (2020); Bentsen et al. (2019).
A wide variety of tests have been developed to diagnose and classify quantum chaos, including spectral form factors Kos et al. (2018); Liu (2018); Bertini et al. (2018), level statistics Berry and Tabor (1977); Bohigas et al. (1984), different definitions of complexityRoberts and Yoshida (2017); Balasubramanian et al. (2022), Loschmidt echos Peres (1984); Jalabert and Pastawski (2001); Gorin et al. (2006) and out-of-time ordered correlators (OTOCs)Sekino and Susskind (2008); Maldacena et al. (2016). Out of these many tools, Krylov Complexity — which can be defined for both states and operators — and OTOCs have emerged as the most comprehensive tests and have become the primary norm for diagnosing quantum chaos in many body systems Parker et al. (2019); Rabinovici et al. (2022); Nandy et al. (2024); Baggioli et al. (2024); Chapman et al. (2024); Jha and Roy (2024); Alishahiha et al. (2024); Swingle (2018); Lin and Motrunich (2018); García-Mata et al. (2023); Schuckert and Knap (2019).
@@ I Introduction
A summary of our results can be seen in Fig. 1, where a simple model inspired by holography — the Hyperbolic Ising model — satisfies multiple criteria of quantum chaos and fast scrambling. We evaluate the scrambling time $t_{s}$ and find a logarithmic lightcone, with $t_{s}\sim\log(N)$, where $N$ is the number of spins. The Lyapunov exponents obtained from OTOCs at finite temperatures decay as $a/\beta$, where $\beta=1/T$ is the inverse temperature and $a$ is a constant. Most importantly, the decay of these Lyapunov exponents for a certain set of parameters saturates the Maldacena-Shenker-Stanford (MSS) bound on chaos Maldacena et al. (2016). We also study the Krylov state complexity (K-complexity), which exhibits the expected ramp-peak-plateau behavior Jha and Roy (2024); Alishahiha et al. (2024), the Krylov Entropy (K-entropy) which rises exponentially at early times, and the Krylov operator complexity which shows the expected exponential-linear-saturation behavior Parker et al. (2019); Rabinovici et al. (2022); Nandy et al. (2024); all of these are indicative of fast scrambling and chaos.
@@
II Hyperbolic Ising Model
Our model of interest, the Hyperbolic Ising model is a mixed field Ising (MFI) model formulated on a one-dimensional hyperbolic space. In Fig. 1 we show how one can visualize this model as a line between two boundary points in $AdS_{3}$ and interpret the effects of the background curvature as site-dependent couplings.
@@ II Hyperbolic Ising Model
In an earlier work Asaduzzaman et al. (2024b), we investigated information propagation and scrambling in this model in the infinite temperature limit, and showed that for a suitable subset of the parameters the lightcone obtained from infinite temperature OTOCs is logarithmic such that the saturation time $t_{s}\sim\log(i)$ where $i$ is the site index which implies that $t_{s}\sim\log(N)$ for a system of $N$ spins. This logarithmic spreading can be seen in Fig. 2 where lines show fits to a logarithmic function, and it is one of the tell-tale signs of fast scrambling Sekino and Susskind (2008). Surprisingly, this simple model with local interactions captures this behavior quite easily due to its unique site-dependent couplings arising from the curved background contrary to the usual expectations that require all-to-all or long-range interactions to achieve such behavior.
III Observables for Chaos and Fast Scrambling
Motivated by the observations of logarithmic lightcones in this model, we will provide further evidence that this model is indeed chaotic and can be classified as a fast-scrambler by investigating finite temperature OTOCs and various complexity measures defined using Krylov subspace methods.
III.1 OTOCs at Finite Temperature
OTOCs are one of the most heavily used diagnostic tools for scrambling and chaos. They are obtained from the double commutator of operators $W(t)=e^{iHt}We^{-iHt}$ and $V$ and can be given as follows.
@@ III.1 OTOCs at Finite Temperature
As discussed earlier, we will choose $J/h=1.0$, and $m=0.05$ such that the model is near the chaotic region of the mixed field Ising model and choose $W_{i}$, and $V_{j}$ as $\sigma_{\frac{N+1}{2}}^{z}$ and $\sigma_{1}^{z}$ respectively and obtain the finite temperature states that go into this calculation by purification techniquesFeiguin and White (2005); Alhambra and Cirac (2021) developed for Matrix Product States (MPS) [^2] For tensor network simulations we used the ITensor Library Fishman et al. (2020); details of our simulation procedure can be found in Appendix. B. White and Feiguin (2004); White (1993); Schollwöck (2005); Vidal (2003); Verstraete et al. (2004); Vidal (2004).
After the desired thermal state is obtained it is then time evolved while measuring $O=1-\tilde{F_{ij}}(t)$ at each time-step. In Fig. 3 we show our results for the OTOCs at $l_{\rm max}=4.0$ at inverse temperatures ranging from $\beta=[0.5,...6.0]$. To extract the Lyapunov exponents we fit $\log(\frac{dO}{dt})$ to $f(x)=at+b$; this approach yields better fits with better $R^{2}$ results compared to exponential fits of the original function.
Our results presented in Fig. 3 (a) confirm the expected exponential rise of the OTOCs with the two dashed lines indicating the fitting range. This range is chosen such that for all the values of $\beta$ the linear fits result in a $R^{2}\geq 0.96$.
Fig. 3 (b) shows the decay of the $\lambda_{\beta}$ as $a/\beta$ where $a$ is some constant. Most importantly we clearly see that as the background curvature is increased the Lyapunov exponents start saturating the Maldacena-Shenker-Stanford (MSS) bound on chaos which is the most famous criterion for fast scrambling.
Figure 4: a) $C_{K}$ using states for $N=11,J/h=1.0,m=0.05$ at different $l_{\rm max}$ values showing the ramp-peak-plateau behavior. b) $S_{K}$ for the same set of parameters showing the expected exponential behavior. c) Lanczos coefficients as a function of the state index $n$ which shows the expected linear dependence on $n$.
Although our results show agreement with the MSS bound we note that this result is highly dependent on choosing the right fitting window. The dependence of $1/\beta$ is less sensitive in changes to this window but the coefficient of the fit is and thus can affect the temperature range where the MSS bound is satisfied. We also note that due to the Trotterized evolution used in both obtaining the thermal state and the time-evolved states, we are limited to moderate values of $\beta$ that can be simulated reliably due to the accumulation of Trotter errors.
@@
III.2 Krylov Observables
Now we switch gears and discuss our observables obtained using Krylov subspace methods. The main advantage of these observables is that they do not require a delicate fitting scheme like OTOCs do to diagnose chaos.
@@ III.2 Krylov Observables
This brings us to our first set of observables that signals a chaotic system in the context of Krylov methods. We start the Lanczos algorithm with the state $\bigotimes_{i=1}^{N}\ket{+^{y}}$, where |$\ket{+^{y}}=\frac{1}{\sqrt{2}}\left(\ket{0}^{y}+i\ket{1}^{y}\right)$. We chose this initial seed choice because in the flat Ising model this state exhibited a clear peak in $C_{K}(t)$ in the chaotic regime of the model Alishahiha et al. (2024). [^3] The choice of the initial seed is an essential part of this algorithm and it can have an effect on the results obtained for $C_{K}(t)$ and $S_{K}(t)$. To alleviate this dependence on the initial seed new algorithms that sample a wide array of initial seeds have been proposed Craps et al. (2025). Using this initial choice we carry out the subspace generation and time evolution and investigate $C_{K}(t)$ and $S_{K}(t)$.
There are three key indicators of quantum chaos obtained using states. First, $C_{K}(t)$ is expected to exhibit ramp-peak-plateau behavior in time. Next, $S_{K}(t)$ should exhibit an exponential behavior at early timescales. Finally, the Lanczos coefficients $b_{n}$ should have a linear relationship with $n$, where $10\leq n\leq\mathbb{K}-$ 1 Parker et al. (2019); Rabinovici et al. (2022); Nandy et al. (2024); Baggioli et al. (2024); Chapman et al. (2024); Jha and Roy (2024).
Our results confirm the expectations described above and show the system becomes more chaotic as the background curvature increases. In Fig. 4 (a) we show the $C_{K}(t)$ where the ramp-peak-plateau behavior is clearly visible and the peak becomes more prominent as we increase $l_{\rm max}$ in line with the expectation that the model becomes more chaotic as $l_{\rm max}$ is increased. In Fig. 4 (b) we show $S_{K}(t)$ which shows the expected exponential rise, the lines show fits to a function of form $f(t)=a(e^{c}-d)$ and the factor $c$ increases with $l_{\rm max}$. Finally, in Fig. 4 (c), we show the Lanczos coefficient $b_{n}$ vs $n$ which shows the linear dependence on $n$, providing us with three more tests signaling chaos.
Figure 5: Operator Complexity for $N=6,J/h=1.0,m=0.05$,. Inset shows the change from exponential to linear behavior around the scrambling times.
@@
The $C_{K}(t)$ obtained using the operators is closely related to operator growth in the full model and hence can be used to identify scrambling behavior and chaos Nandy et al. (2024). For a fast scrambler, operator complexity is expected to grow exponentially at early timescales followed by a linear increase and finally, they saturate/oscillate about some fixed average value related to the dimension of the Krylov subspace Barbón et al. (2019).
Fig. 5 confirms this expected behavior for a chaotic system with the inset showing the early time dynamics of the complexity which is fitted to an exponential followed by a linear function. Confirming our discovery of scrambling in the Hyperbolic Ising model.
IV Conclusions and Outlook
In this paper, we investigated chaos and scrambling in the Hyperbolic Ising model which is a nearest neighbor mixed field Ising model with site-dependent couplings using tensor network and Krylov subspace methods and showed that the observables obtained from both of these independent techniques show signatures of fast scrambling and chaos. More specifically we had three main observables with such a signature.
@@ IV Conclusions and Outlook
Acknowledgments
We thank Raghav Govind Jha and Bharath Sambasivam for valuable discussions. AFK and GCT acknowledge financial support from the National Science Foundation under award No. 2325080: PIF: Software-Tailored Architecture for Quantum Co-Design. AS is supported by U.S. Department of Energy grant DE-SC0019139. We acknowledge the computing resources provided by North Carolina State University High Performance Computing Services Core Facility (RRID:${\rm SCR\_}022168$).
@@ Appendix A Krylov Subspace Methods
To obtain the time evolution of these vectors, recursion relations in Eq. LABEL:eq:rec_st8 can be written in a vector-matrix linear ordinary differential equation form given in Eq. 9.

### v2

# Fast Scrambling in the Hyperbolic Ising Model

@@ Abstract
Introduction— Quantum many-body chaos and scrambling have attracted a great deal of attention in recent years, from more theoretical endeavors like studying the scrambling dynamics of black holes using their dual descriptions Kobrin et al. (2021); Fu and Sachdev (2016); Asaduzzaman et al. (2024a); Brown et al. (2023); Schuster et al. (2022a); Cotler et al. (2017); Joshi et al. (2022) to more practical applications like investigating scrambling in chemical reactions to quantum cryptography Zhang et al. (2024); Chamon et al. (2024) and many moreLandsman et al. (2019); Schuster et al. (2022b); Blok et al. (2021).
In almost all of these studies, the Sachdev-Ye-Kitaev (SYK) model has been the center of attention as the quintessential example of a model that exhibits fast scrambling and quantum chaos Sachdev and Ye (1993); noa (a, b); Maldacena and Stanford (2016). However, due to the all-to-all and random interactions in this model, any kind of classical and quantum simulation can be very challenging and often requires various truncations and approximations Xu et al. (2020); García-García et al. (2021); Cao et al. (2020); Sünderhauf et al. (2019). In this letter, we advocate for a new model that is also maximally chaotic and exhibits fast scrambling but with only local site-dependent interactions.  [^1] There are other examples of fast scrambling models besides the SYK model but to the best of our knowledge they all require longer-range interactions than nearest neighbor interactions to be present in the model Hashizume et al. (2021); Li et al. (2020); Lashkari et al. (2013); Belyansky et al. (2020); Yin and Lucas (2020); Bentsen et al. (2019); Chen and Zhou (2019).
A wide variety of tests have been developed to diagnose and classify quantum chaos,
including spectral form factors Kos et al. (2018); Liu (2018); Bertini et al. (2018), level statistics Berry and Tabor (1977); Bohigas et al. (1984), different definitions of complexityRoberts and Yoshida (2017); Balasubramanian et al. (2022), Loschmidt echos Peres (1984); Jalabert and Pastawski (2001); Gorin et al. (2006) and out-of-time ordered correlators (OTOCs)Sekino and Susskind (2008); Maldacena et al. (2016); Swingle et al. (2016); Xu and Swingle (2020); Swingle (2018a). Out of these many tools, Krylov Complexity — which can be defined for both states and operators — and OTOCs have emerged as the most comprehensive tests and have become the primary norm for diagnosing quantum chaos in many body systems Parker et al. (2019); Rabinovici et al. (2022); Nandy et al. (2024); Baggioli et al. (2024); Chapman et al. (2024); Jha and Roy (2024); Alishahiha et al. (2024); Swingle (2018b); Lin and Motrunich (2018); García-Mata et al. (2023); Schuckert and Knap (2019); Balasubramanian et al. (2025).
@@
A summary of our results can be seen in Fig. 1, where a simple model inspired by holography — the Hyperbolic Ising model — satisfies multiple criteria of quantum chaos and fast scrambling. We evaluate the scrambling time $t_{s}$ and find a logarithmic lightcone, with $t_{s}\sim\log(N)$, where $N$ is the number of spins. The Lyapunov exponents obtained from OTOCs at finite temperatures decay as $a/\beta$, where $\beta=1/T$ is the inverse temperature and $a$ is a constant. Most importantly, the decay of these Lyapunov exponents for a certain set of parameters saturates the Maldacena-Shenker-Stanford (MSS) bound on chaos Maldacena et al. (2016). We also study the Krylov state complexity (K-complexity), which exhibits the expected ramp-peak-plateau behavior Jha and Roy (2024); Alishahiha et al. (2024), the Krylov Entropy (K-entropy), which rises exponentially at early times, and the Krylov operator complexity, which shows the expected exponential-linear-saturation behavior Parker et al. (2019); Rabinovici et al. (2022); Nandy et al. (2024); all of these are indicative of fast scrambling and chaos.
@@
Hyperbolic Ising Model— Our model of interest, the Hyperbolic Ising model, is a mixed field Ising (MFI) model formulated on a one-dimensional hyperbolic space. In Fig. 1, we show how one can visualize this model as a line between two boundary points in $AdS_{3}$ and interpret the effects of the background curvature as site-dependent couplings.
@@
In an earlier work Asaduzzaman et al. (2024b), we investigated information propagation and scrambling in this model in the infinite temperature limit, and showed that for a suitable subset of the parameters the lightcone obtained from infinite temperature OTOCs is logarithmic such that the saturation time $t_{s}\sim\log(i)$ where $i$ is the site index which implies that $t_{s}\sim\log(N)$ for a system of $N$ spins. This logarithmic spreading can be seen in Fig. 2, where lines show fits to a logarithmic function, and it is one of the tell-tale signs of fast scrambling Sekino and Susskind (2008). Surprisingly, this simple model with local interactions captures this behavior quite easily due to its unique site-dependent couplings arising from the curved background, contrary to the usual expectations that require all-to-all or long-range interactions to achieve such behavior.
Observables for Chaos and Fast Scrambling— Motivated by the observations of logarithmic lightcones in this model, we will provide further evidence that this model is indeed chaotic and can be classified as a fast-scrambler by investigating finite temperature OTOCs and various complexity measures defined using Krylov subspace methods.
OTOCs at Finite Temperature— OTOCs are one of the most heavily used diagnostic tools for scrambling and chaos. They are obtained from the double commutator of operators $W(t)=e^{iHt}We^{-iHt}$ and $V$ and can be given as follows.
@@
As discussed earlier, we will choose $J/h=1.0$, and $m=0.05$ such that the model is near the chaotic region of the mixed field Ising model and choose $W_{i}$, and $V_{j}$ as $\sigma_{\frac{N+1}{2}}^{z}$ and $\sigma_{1}^{z}$ respectively and obtain the finite temperature states that go into this calculation by purification techniquesFeiguin and White (2005a); Alhambra and Cirac (2021); Feiguin and White (2005b) developed for Matrix Product States (MPS) [^2] For tensor network simulations we used the ITensor Library Fishman et al. (2020); details of our simulation procedure can be found in Appendix. B. White and Feiguin (2004); White (1993); Schollwöck (2005); Vidal (2003); Verstraete et al. (2004); Vidal (2004).
After the desired thermal state is obtained, it is then time evolved while measuring $O=1-\tilde{F_{ij}}(t)$ at each time-step. In Fig. 3 we show our results for the OTOCs at $l_{\rm max}=4.0$ at inverse temperatures ranging from $\beta=[0.5,...6.0]$. To extract the Lyapunov exponent,s we fit $\log(\frac{dO}{dt})$ to $f(x)=at+b$; this approach yields better fits with better $R^{2}$ results compared to exponential fits of the original function.
Our results presented in Fig. 3 (a) confirm the expected exponential rise of the OTOCs with the two dashed lines indicating the fitting range. This range is chosen such that for all the values of $\beta$, the linear fits result in a $R^{2}\geq 0.96$.
Fig. 3 (b) shows the decay of the $\lambda_{\beta}$ as $a/\beta$ where $a$ is some constant. Most importantly, we clearly see that as the background curvature is increased, the Lyapunov exponents start saturating the Maldacena-Shenker-Stanford (MSS) bound on chaos, which is the most famous criterion for fast scrambling.
Figure 4: a) $C_{K}$ using states for $N=11,J/h=1.0,m=0.05$ at different $l_{\rm max}$ values showing the ramp-peak-plateau behavior. b) $S_{K}$ for the same set of parameters showing the expected exponential behavior. c) Lanczos coefficients $b_{n}$ as a function of the state index $n$, which displays the Lanczos descent.
Although our results show agreement with the MSS bound, we note that this result is highly dependent on choosing the right fitting window. The dependence of $1/\beta$ is less sensitive in changes to this window,w but the coefficient of the fit is and thus can affect the temperature range where the MSS bound is satisfied. We also note that due to the Trotterized evolution used in both obtaining the thermal state and the time-evolved states, we are limited to moderate values of $\beta$ that can be simulated reliably due to the accumulation of Trotter errors.
@@
Krylov Observables— Now we switch gears and discuss our observables obtained using Krylov subspace methods. The main advantage of these observables is that they do not require a delicate fitting scheme like OTOCs do to diagnose chaos.
@@
This brings us to our first set of observables that signal a chaotic system in the context of Krylov methods. We start the Lanczos algorithm with the state $\bigotimes_{i=1}^{N}\ket{+^{y}}$, where |$\ket{+^{y}}=\frac{1}{\sqrt{2}}\left(\ket{0}^{y}+i\ket{1}^{y}\right)$. We chose this initial seed choice because in the flat Ising model, this state exhibited a clear peak in $C_{K}(t)$ in the chaotic regime of the model Alishahiha et al. (2024). [^3] The choice of the initial seed is an essential part of this algorithm, and it can have an effect on the results obtained for $C_{K}(t)$ and $S_{K}(t)$. To alleviate this dependence on the initial seed new algorithms that sample a wide array of initial seeds have been proposed Craps et al. (2025). Using this initial choice, we carry out the subspace generation and time evolution and investigate $C_{K}(t)$ and $S_{K}(t)$.
There are two key indicators of quantum chaos obtained using states. First, $C_{K}(t)$ is expected to exhibit ramp-peak-plateau behavior in time. Next, $S_{K}(t)$ should exhibit an exponential behavior at early timescales. Parker et al. (2019); Rabinovici et al. (2022); Nandy et al. (2024); Baggioli et al. (2024); Chapman et al. (2024); Jha and Roy (2024).
Our results confirm the expectations described above and show that the system becomes more chaotic as the background curvature increases. In Fig. 4 (a) we show the $C_{K}(t)$ where the ramp-peak-plateau behavior is clearly visible and the peak becomes more prominent as we increase $l_{\rm max}$ in line with the expectation that the model becomes more chaotic as $l_{\rm max}$ is increased. In Fig. 4 (b) we show $S_{K}(t)$ which shows the expected exponential rise, the lines show fits to a function of form $f(t)=a(e^{c}-d)$ and the factor $c$ increases with $l_{\rm max}$.
Figure 5: Operator Complexity for $N=6,J/h=1.0,m=0.05$,. The inset shows the change from exponential to linear behavior around the scrambling times.
@@
The $C_{K}(t)$ obtained using the operators is closely related to operator growth in the full model and hence can be used to identify scrambling behavior and chaos Nandy et al. (2024). For a fast scrambler, operator complexity is expected to grow exponentially at early timescales followed by a linear increase and finally, they saturate/oscillate about some fixed average value related to the dimension of the Krylov subspace. And the Lanczos coefficients $b_{n}$ should have a linear relationship with $n$, where $10\leq n\leq\mathbb{K}-$ 1. This shape that the Lanczos coefficients display is generally referred to as the Lanczos descent Barbón et al. (2019).
In Fig. 4 (c), we show the Lanczos coefficient $b_{n}$ vs $n$, which shows the Lanczos decent behavior, and Fig. 5 confirms this expected behavior for a chaotic system, with the inset showing the early time dynamics of the complexity, which is fitted to an exponential function followed by a linear function. Confirming our discovery of scrambling in the Hyperbolic Ising model.
Conclusions and Outlook— In this paper, we investigated chaos and scrambling in the Hyperbolic Ising model which is a nearest neighbor mixed field Ising model with site-dependent couplings using tensor network and Krylov subspace methods and showed that the observables obtained from both of these independent techniques show signatures of fast scrambling and chaos. More specifically we had three main observables with such a signature.
@@
Acknowledgments— We thank Raghav Govind Jha and Bharath Sambasivam for valuable discussions. AFK and GCT acknowledge financial support from the National Science Foundation under award No. 2325080: PIF: Software-Tailored Architecture for Quantum Co-Design. AS is supported by U.S. Department of Energy grant DE-SC0019139. We acknowledge the computing resources provided by North Carolina State University High Performance Computing Services Core Facility (RRID:${\rm SCR\_}022168$).
@@ Appendix A Krylov Subspace Methods
To obtain the time evolution of these vectors, recursion relations in Eq. 8 can be written in a vector-matrix linear ordinary differential equation form given in Eq. 9.